Which statement is true for the single price monopoly that is operating on the demand curve P=75-5Q at a point where price is R37,50
1) the firm should definitely not lower its price
2) the firm should definitely lower its price
3) if the marginal cost is 0, the firm should definitely increase its price
4) None of the above
1
Expert's answer
2020-05-13T11:09:50-0400
"p=75-5Q"


"TR=pQ"


"TR=75Q-5Q^2"


"MR=75-10Q"

"75-10Q=0"


"Q=7.5"


"p=75-5\times7.5=37.5"

Answer:1